40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 2598     Entry time: Fri Jul 2 11:27:01 2021
Author: Radhika 
Type: DailyProgress 
Category: Cryo vacuum chamber 
Subject: Ccooling updates 6/29-7/2 

[Radhika, Stephen]

On Tuesday 6/29, we opened up Megastat after a weekend of warmup. Our goals were to 1) check on the RTD attached(?) to the cold plate to confirm that it had de-bonded, and 2) to check on the foil and mylar inside the chamber and around the copper braid. 

Indeed the tip of the cold head RTD had detached [attachment 1], but the mylar "cap" I created for insulation held up well. Inside the chamber, we noticed some shorting between aluminum foil on the inside of the outer lid and the outside of the inner lid [attachment 2]. Also, there was a gap in mylar around the copper braid, so some of the braid was visible. We also noticed that the strip of peek sheet inside the opening of the inner shield had slipped out of place, exposing the copper braid to the aluminum [attachment 8]. 

We rebonded the loose RTD to the cold head [attachment 3] and re-inserted the mylar "cap" [attachment 4 and 10]. Stephen suggested we add an additional piece of mylar to cover the portion of the cold head directly above the braid [attachment 5]. 

To prevent shorting inside the chamber, we trimmed the excess aluminum foil wrapped around the non-coated sides of the inner and outer shields [attachment 6]. We adjusted the mylar sheets around the copper braid to cover the previous gap, and used a mirror to ensure the bottom of the braid was covered [attachment 7]. We also realigned the peek sheet strip with the opening of the inner shield [attachment 9]. 

We reattached the heater to the workpiece [attachment 11] and the workpiece RTD with the spring clamp [attachment 12]. We placed back the inner and outer shield lids [attachment 13-14], closed up the chamber, and pumped down. Since I have been working on remotely extracting temperature data from the CTC-100, we chose to wait until the script was ready before turning on the cryocooler.

On Thursday 7/1, I got the temperature extraction script ready and turned on the cryocooler at ~5:15 PM. Fun addition: attachment 15 shows successful communication to the CTC-100 (I sent a command to display "hello" on the screen :D)

On Friday 7/2, I plotted the temperature trends so far [attachment 16]. As of 12pm Friday, the workpiece is at 177K. We will let the chamber cool over the weekend and see where we are on Tuesday.


Attachment 1: IMG_9126.jpeg  174 kB  Uploaded Fri Jul 2 12:46:27 2021  | Hide | Hide all
Attachment 2: IMG_9136.jpeg  197 kB  Uploaded Fri Jul 2 12:46:54 2021  | Hide | Hide all
Attachment 3: IMG_9147.jpeg  84 kB  Uploaded Fri Jul 2 12:48:32 2021  | Hide | Hide all
Attachment 4: IMG_9148.jpeg  209 kB  Uploaded Fri Jul 2 12:48:43 2021  | Hide | Hide all
Attachment 5: IMG_9145.jpeg  182 kB  Uploaded Fri Jul 2 12:48:51 2021  | Hide | Hide all
Attachment 6: IMG_9144.jpeg  268 kB  Uploaded Fri Jul 2 12:51:27 2021  | Hide | Hide all
Attachment 7: IMG_9142.jpeg  153 kB  Uploaded Fri Jul 2 12:51:56 2021  | Hide | Hide all
Attachment 8: IMG_9139.jpeg  234 kB  Uploaded Fri Jul 2 12:55:32 2021  | Hide | Hide all
Attachment 9: IMG_9152.jpeg  351 kB  Uploaded Fri Jul 2 12:56:28 2021  | Hide | Hide all
Attachment 10: IMG_9149.jpeg  138 kB  Uploaded Fri Jul 2 12:57:22 2021  | Hide | Hide all
Attachment 11: IMG_9153.jpeg  210 kB  Uploaded Fri Jul 2 12:59:02 2021  | Hide | Hide all
Attachment 12: IMG_9151.jpeg  409 kB  Uploaded Fri Jul 2 12:59:13 2021  | Hide | Hide all
Attachment 13: IMG_9154.jpeg  509 kB  Uploaded Fri Jul 2 13:00:50 2021  | Hide | Hide all
Attachment 14: IMG_9156.jpeg  396 kB  Uploaded Fri Jul 2 13:01:00 2021  | Hide | Hide all
Attachment 15: IMG_9158.jpeg  224 kB  Uploaded Fri Jul 2 13:04:29 2021  | Hide | Hide all
Attachment 16: cooldown_7_1-7_2.pdf  14 kB  Uploaded Fri Jul 2 13:07:22 2021  | Hide | Hide all
ELOG V3.1.3-